A copy company wants to expand production. It currently has 20 workers who share eight copiers.Two months ago, the firm added two copiers,and output increased by 100,000 pages per day.One month ago, they added five workers, and productivity also increased by 50,000 pages per day. Copiers cost about twice as much as workers. Would you recommend they hire another employee or buy another copier? (OH BTW PLEASE MAKE SURE THIS IS NOT A REPEAT ANSWER FROM ONE OF THE OTHER ANSWERS ALREADY POSTED)
Answer #1

By adding two copiers amount increased = 100,000 pages

By adding 5 workers, total workers = 25

Increase in pages = Increase in productivity * number of workers

= 50,000*25 =1,250,000 pages

Cost of additional copiers be 2x

Cost of two copier = 2*2x = 4x

Pages increased per x = 100000/4 = 25000

Therefore cost of worker would be half of copier which is x

Cost of five workers = 5x

Pages increased per x = 1,250,000/5 = 500,000

Because, per unit cost the pages increased by adding worker is higher than that of adding a copier it is advisable that additional workers are hired in this regard which would be better than buying a copier